# On Structural Properties of Risk-Averse Optimal Stopping Problems

### v2

@@ 4 Characterize the structure of the optimal policy
  • Continuation is optimal at time $t$ if and only if (iff) $L_{t}(x_{t})\leq 0$.
  • For each $t$ and some fixed $i$, if $L_{t}(x_{t})$ is decreasing (resp., increasing) in $x_{t,i}$ for every fixed $x_{t,-i}$, then a control limit optimal policy in the $i$-th dimension exists: there exists a threshold function $\overline{x}_{t}^{i}:{\mathbb{R}}^{n-1}\mapsto{\mathbb{R}}\cup\{\pm\infty\}$ such that it is optimal to continue (resp., stop) at time $t$ iff $x_{t,i}\geq\overline{x}_{t}^{i}(x_{t,-i})$.
@@
Proposition 4.12 indicates that in the risk-neutral case (each $\rho_{t}$ is a conditional expectation), if $Q_{t}(\cdot|x)$ is stochastically increasing in $x\in\mathcal{X}$, the condition that $M_{t}$ is decreasing suffices for the existence of control limit optimal policies. Moreover, this condition is preferable to the classical modularity conditions (C1) and (C2) because:
  • It is weaker. For example, consider a time-homogeneous MDP, so we can drop time subscripts on the cost functions and transition law. Then (C1) implies that $\rho_{t}\big(v_{T}\big(\tilde{X}(x)\big)\big)=\rho_{t}\big(s\big(\tilde{X}(x)\big)\big)$ is decreasing in $x$, together with (C2), implies that $M_{t}$ is decreasing.
  • It is easier to check, since $M_{t}$ depends only on primitives (cost functions, the transition law, and the one-step risk mappings), whereas establishing (C1) is more challenging.
However, since $\rho_{t}$ is subadditive, in general, $L_{t}$ no longer inherits the monotonicity of $M_{t}$. Guided by Proposition 4.12 and the analysis that follows, we therefore seek conditions in risk-averse settings under which the structure of $M_{t}$ still certifies the optimal policy structure. In Section 4.1, we establish that if both the one-step risk mappings and the state vector satisfy appropriate comonotonicity conditions, subadditivity tightens to additivity, restoring the risk-neutral argument. In Section 4.2, we give conditions under which the sign of $L_{t}$ is determined entirely by $M_{t}$; consequently, a one-step look-ahead policy is optimal.
@@ 4.1 Comonotone Risk Measures and State Vectors
  • Random variables $X,Y$ on measurable space $(\Omega,\mathcal{F})$ are comonotone if $(X(\omega)-X(\omega^{\prime}))(Y(\omega)-Y(\omega^{\prime}))\geq 0$ for all $\omega,\omega^{\prime}\in\Omega$.
@@
The random vector $\tilde{X}_{t+1}(x_{t})$ is comonotone for each $t$ and $x_{t}\in\mathcal{X}$.
@@
Since $L_{t+1}$ is decreasing, $\min\{0,L_{t+1}(\cdot)\}$ is also decreasing. Fix $x_{t}\in\mathcal{X}$. By Assumption 7 and Proposition 4.15, there exists a random variable $W_{t}$ (possibly depending on $x_{t}$) and a mapping $f:{\mathbb{R}}\mapsto\mathcal{X}$ with each component $f_{i}:{\mathbb{R}}\mapsto{\mathbb{R}}$ increasing such that $\tilde{X}_{t+1}(x_{t})=f(W_{t})$. Hence the compositions $s_{t+1}(f(\cdot)),\min\{0,L_{t+1}(f(\cdot))\}$ are decreasing. By Proposition 4.15, $s_{t+1}\big(\tilde{X}_{t+1}(x_{t})\big)$ and $\min\big\{0,L_{t+1}\big(\tilde{X}_{t+1}(x_{t})\big)\big\}$ are comonotone. Since $\rho_{t}$ is comonotone, (6) yields
@@ 4.2 One-Step Look-Ahead Optimal Policy
The one-step look-ahead policy is determined by the sign of the one-step loss $M_{t}$ as follows: for each $t$,
$d_{t}(x_{t})=\begin{cases}S&M_{t}(x_{t})\geq 0,\\ C&M_{t}(x_{t})<0,\end{cases}$
i.e., stopping at time $t$ whenever deferring termination to $t+1$ increases risk. Because this policy is far simpler to compute than solving the DP, it is desirable to know when it is optimal. Proposition 4.12 states that the optimal action is characterized by the sign of $L_{t}$, so the one-step look-ahead policy is optimal iff $M_{t}$ and $L_{t}$ have the same sign. We now give a sufficient condition.
@@
For every $t$ and $x_{t}\in\mathcal{X}$, if $M_{t}(x_{t})\geq 0$, then $M_{t+1}\big(\tilde{X}_{t+1}(x_{t})\big)\geq 0$ a.s.
@@
Under Assumptions 1 and 9, $M_{t}$ and $L_{t}$ have the same sign for all $t$. Consequently, the one-step look-ahead policy is optimal.
Proof 4.23.
@@
Conversely, by the second inequality in 5, $L_{t}(x_{t})\leq M_{t}(x_{t})$ for all $x_{t}$. Therefore, $M_{t}(x_{t})\leq 0\implies L_{t}(x_{t})\leq 0$. Thus, $M_{t}$ and $L_{t}$ have the same sign, and the one-step look-ahead policy is optimal at time $t$.
Consider the componentwise partial order $\preceq$ on $\mathcal{X}\subseteq{\mathbb{R}}^{n}$. By Theorem 4.22, if, in addition, $M_{t}$ is monotone for every $t$, then the one-step look-ahead policy is optimal and of control limit-type. We next present verifiable conditions that ensure Assumption 9 given $M_{t}$ is monotone.
@@
For each $t$: (i) $M_{t}(x)$ is increasing (resp., decreasing) in $x$; (ii) for all $x$, $M_{t}(x)$ and $M_{t+1}(x)$ have the same sign; and (iii) $x\preceq\tilde{X}_{t+1}(x)$ (resp., $\succeq$) a.s.
Condition (ii) is automatic in time-invariant systems. Condition (iii) captures “non-improving” dynamics. For instance, in transplant decision-making or machine maintenance models where larger state $X_{t}$ denotes poorer health, $x\preceq\tilde{X}_{t+1}(x)$ a.s. indicates that, upon continuation, the state almost surely does not improve.
@@
Under Assumptions 1 and 10, the one-step look-ahead policy is optimal and is a control limit policy.
Proof 4.25.
Consider the case $M_{t}$ is increasing (the decreasing case is analogous with reversed inequalities). Fix $x$ with $M_{t}(x)\geq 0$. By (i) and (iii) in Assumption 10, $M_{t}\big(\tilde{X}_{t+1}(x)\big)\geq 0$ a.s. By (ii), $M_{t+1}\big(\tilde{X}_{t+1}(x)\big)\geq 0$ a.s., so Theorem 4.22 implies the one-step look-ahead policy is optimal. Finally, the monotonicity of $M_{t}(x)$ implies a control limit structure (as in Lemma 4.11).
@@
It is straightforward to verify that Assumption 10 holds. Hence, by Corollary 4.24, the one-step look-ahead policy is optimal and of control limit-type. $\square$
@@
In general, since $L_{t}(x_{t})\leq M_{t}(x_{t})$ for all $x_{t}\in\mathcal{X}$, $M_{t}(x_{t})\leq 0$ guarantees that continuation is optimal, whereas $M_{t}(x_{t})\geq 0$ is inconclusive. If, for each $t$, $M_{t}$ is increasing (resp., decreasing) in its $i$-th coordinate, then for any fixed $x_{t,-i}$, there exists a threshold $\overline{x}_{t}^{i}(x_{t,-i})\in{\mathbb{R}}$ such that whenever $x_{t,i}\leq\overline{x}_{t}^{i}(x_{t,-i})$ (resp., $\geq$), $M_{t}(x_{t})\leq 0$, and thus continuation is optimal. On the complementary side of the threshold, $M_{t}$ does not determine the optimal action. This resembles the $(s,S,A,p)$ policy structure in [32].
@@ 5 Monotonicity of the optimal control limits
  • Monotonicity in time. Consider a time-homogeneous model (the transition kernel and costs functions are fixed, and the one-step risk mappings have the same functional form across time). If $L_{t}(x)$ is decreasing in $x_{i}$ for each $t$ and some fixed $i$, then a control limit optimal policy in the $i$-th dimension exists, and the optimal control limit $\overline{x}_{t,i}(x_{t,-i})$ is increasing in $t$ for any fixed $x_{t,-i}$.
  • Monotonicity in risk-aversion level. Consider two time-homogeneous risk-averse optimal stopping instances with identical transition kernels and cost functions, and one-step risk mappings $\{\rho_{t}^{1}\}$ and $\{\rho_{t}^{2}\}$, respectively. Let the corresponding value functions be $v_{t}^{1}$ and $v_{t}^{2}$. For some fixed $i$, suppose a control limit optimal policy in the $i$-th dimension exists in both instances, with control limit functions denoted by $\overline{x}_{t,i}^{1}$ and $\overline{x}_{t,i}^{2}$, respectively. If the first instance is more risk-averse, i.e., $\rho_{t}^{1}(Z)\geq\rho_{t}^{2}(Z)$ for every $t$ and every $Z\in\mathcal{Z}_{t+1}$, then $v_{t}^{1}\geq v_{t}^{2}$ and $\overline{x}_{t,i}^{1}(x_{t,-i})\geq\overline{x}_{t,i}^{2}(x_{t,-i})$ for every $t$ and $x_{t,-i}$.
@@
The third statement formalizes the effect of increased risk aversion under identical dynamics and costs. Its proof is a straightforward backward-induction argument and is omitted. For illustration, consider one-step mean–CVaR mappings $\rho_{t}^{i}(\cdot)=(1-\alpha_{i}){\mathbb{E}}(\cdot|\mathcal{F}_{t})+\alpha_{i}\operatorname*{CVaR}_{\gamma,t}(\cdot)$ (with fixed $\gamma\in[0,1]$) or pure CVaR mappings $\rho_{t}^{i}(\cdot)=\operatorname*{CVaR}_{\alpha_{i},t}(\cdot)$ for $i=1,2$. If $0\leq\alpha_{2}\leq\alpha_{1}\leq 1$, then $\rho_{t}^{1}(Z)\geq\rho_{t}^{2}(Z)$ for every $Z\in\mathcal{Z}_{t+1}$. Intuitively, $\rho_{t}^{1}$ places more weight on the tail or on a more extreme tail and is therefore more risk-averse. Consequently, $v_{t}^{1}\geq v_{t}^{2}$, and the more risk-averse instance tends to stop earlier.

### v3

@@ 4 Characterization of the structure of the optimal policy
  • Continuation $C$ is optimal at time $t$ if and only if (iff) $L_{t}(x_{t})\leq 0$.
  • For each $t$ and some fixed $i$, if $L_{t}(x_{t})$ is decreasing (resp., increasing) in $x_{t,i}$ for each fixed $x_{t,-i}$, a control limit optimal policy in the $i$-th dimension exists: there exists a threshold function $\overline{x}_{t}^{i}:{\mathbb{R}}^{n-1}\mapsto{\mathbb{R}}\cup\{\pm\infty\}$ such that it is optimal to continue (resp., stop) at time $t$ iff $x_{t,i}\geq\overline{x}_{t}^{i}(x_{t,-i})$.
@@
Proposition 4.12 indicates that in the risk-neutral case (each $\rho_{t}$ is a conditional expectation), if $Q_{t}(\cdot|x)$ is stochastically increasing in $x\in\mathcal{X}$, the condition that $M_{t}$ is decreasing suffices for the existence of control limit optimal policies. Moreover, this condition is preferable to the classical modularity conditions (C1) and (C2). First, it is easier to check, since $M_{t}$ depends only on primitives (cost functions, the transition law, and the one-step risk mappings), whereas establishing (C1) is more challenging. Second, it is weaker, as illustrated below.
Example 4.14.
Consider a time-homogeneous risk-neutral MDP where each $\rho_{t}$ is a conditional expectation, so we can drop time subscripts on the cost functions, risk functionals, and transition law. Then (C1) implies that $\rho\big(s\big(\tilde{X}(x)\big)\big)=\rho\big(v_{T}\big(\tilde{X}(x)\big)\big)$ is decreasing in $x$, which, together with (C2), implies that $M_{t}$ is decreasing. $\square$
However, since $\rho_{t}$ is subadditive, in the general risk-averse case, $L_{t}$ no longer inherits the monotonicity of $M_{t}$. Guided by Proposition 4.12 and the analysis that follows, we therefore seek conditions in risk-averse settings under which the structure of $M_{t}$ still implies an optimal policy structure. In Section 4.1, we establish that if both the one-step risk mappings and the state vector satisfy appropriate comonotonicity conditions, subadditivity tightens to additivity, restoring the risk-neutral argument. In Section 4.2, we give conditions under which the sign of $L_{t}$ is determined entirely by $M_{t}$; consequently, a one-step look-ahead policy is optimal.
@@ 4.1 Comonotone Risk Measures and State Vectors
  • Random variables $X,Y$ on measurable space $(\Omega,\mathcal{F})$ are comonotone if $(X(\omega)-X(\omega^{\prime}))(Y(\omega)-Y(\omega^{\prime}))\geq 0,~\forall\omega,\omega^{\prime}\in\Omega$.
@@
The random vector $\tilde{X}_{t+1}(x_{t})$ is comonotone, $\forall t$ and $x_{t}\in\mathcal{X}$.
@@
Since $L_{t+1}$ is decreasing, $\min\{0,L_{t+1}(\cdot)\}$ is also decreasing. Fix $x_{t}\in\mathcal{X}$. By Assumption 7 and Proposition 4.16, there exists a random variable $Y_{t}$ and a mapping $f:{\mathbb{R}}\mapsto\mathcal{X}$ with each component $f_{i}:{\mathbb{R}}\mapsto{\mathbb{R}}$ increasing such that $\tilde{X}_{t+1}(x_{t})=f(Y_{t})$. Hence the compositions $s_{t+1}(f(\cdot)),\min\{0,L_{t+1}(f(\cdot))\}$ are decreasing. By Proposition 4.16, $s_{t+1}\big(\tilde{X}_{t+1}(x_{t})\big)$ and $\min\big\{0,L_{t+1}\big(\tilde{X}_{t+1}(x_{t})\big)\big\}$ are comonotone. Since $\rho_{t}$ is comonotone, (6) yields
@@ 4.2 One-Step Look-Ahead Optimal Policy
We define a one-step look-ahead policy $\overline{d}=\{\overline{d}_{0},\cdots,\overline{d}_{T-1}\}$ as follows. For each $t$, the action is determined by the sign of the one-step loss $M_{t}$:
$\overline{d}_{t}(x_{t})=\begin{cases}S&M_{t}(x_{t})\geq 0,\\ C&M_{t}(x_{t})<0,\end{cases}$
i.e., stopping at time $t$ whenever deferring termination to $t+1$ increases risk. Because this policy is far simpler to compute than solving the DP, it is desirable to know when it is optimal. Proposition 4.12 states that the optimal action is characterized by the sign of $L_{t}$, so the one-step look-ahead policy $\overline{d}$ is optimal iff $M_{t}$ and $L_{t}$ have the same sign. We now give a sufficient condition.
@@
for each $t$ and $x_{t}\in\mathcal{X}$, if $M_{t}(x_{t})\geq 0$, then $M_{t+1}\big(\tilde{X}_{t+1}(x_{t})\big)\geq 0$ a.s.
@@
Under Assumptions 1 and 9, $M_{t}$ and $L_{t}$ have the same sign, $\forall t$. Consequently, the one-step look-ahead policy $\overline{d}$ is optimal.
Proof 4.24.
@@
Conversely, by the second inequality in 5, $L_{t}(x_{t})\leq M_{t}(x_{t}),~\forall x_{t}$. Therefore, $M_{t}(x_{t})\leq 0\implies L_{t}(x_{t})\leq 0$. Thus, $M_{t}$ and $L_{t}$ have the same sign, and the one-step look-ahead policy $\overline{d}$ is optimal at time $t$.
By Theorem 4.23, if, in addition, $M_{t}$ is monotone for each $t$, then the one-step look-ahead policy $\overline{d}$ is optimal and of control limit-type. We next present verifiable sufficient conditions that ensure Assumption 9 given $M_{t}$ is monotone.
@@
For each $t$: (i) $M_{t}(x)$ is increasing (resp., decreasing) in $x$; (ii) $\forall x$, $M_{t}(x)$ and $M_{t+1}(x)$ have the same sign; and (iii) $x\preceq_{c}\tilde{X}_{t+1}(x)$ (resp., $\succeq_{c}$) a.s.
Condition (ii) is automatic in time-invariant systems. Condition (iii) captures “non-improving” dynamics. For instance, in transplant decision-making or machine maintenance models where larger state $X_{t}$ denotes poorer health, $x\preceq_{c}\tilde{X}_{t+1}(x)$ a.s. indicates that, upon continuation, the state almost surely does not improve.
@@
Under Assumptions 1 and 10, the one-step look-ahead policy $\overline{d}$ is optimal and is a control limit policy.
Proof 4.26.
Consider the case $M_{t}$ is increasing (the decreasing case is analogous with reversed inequalities). Fix $x$ with $M_{t}(x)\geq 0$. By (i) and (iii) in Assumption 10, $M_{t}\big(\tilde{X}_{t+1}(x)\big)\geq 0$ a.s. By (ii), $M_{t+1}\big(\tilde{X}_{t+1}(x)\big)\geq 0$ a.s., so Theorem 4.23 implies $\overline{d}$ is optimal. Finally, the monotonicity of $M_{t}(x)$ implies a control limit structure (as in Lemma 4.11).
@@
It is straightforward to verify that Assumption 10 holds. Hence, by Corollary 4.25, $\overline{d}$ is optimal and of control limit-type. $\square$
@@
In general, since $L_{t}(x_{t})\leq M_{t}(x_{t}),~\forall x_{t}\in\mathcal{X}$, $M_{t}(x_{t})\leq 0$ guarantees that continuation is optimal, whereas $M_{t}(x_{t})\geq 0$ is inconclusive. If, for each $t$, $M_{t}$ is increasing (resp., decreasing) in its $i$-th coordinate, then for any fixed $x_{t,-i}$, there exists a threshold $\overline{x}_{t}^{i}(x_{t,-i})\in{\mathbb{R}}$ such that whenever $x_{t,i}\leq\overline{x}_{t}^{i}(x_{t,-i})$ (resp., $\geq$), $M_{t}(x_{t})\leq 0$, and thus continuation is optimal. On the opposite side of the threshold, $M_{t}$ does not determine the optimal action. This resembles the $(s,S,A,p)$ policy structure in [32].
@@ 5 Monotonicity of the optimal control limits
  • Monotonicity in time. Consider a time-homogeneous model (the transition kernel and cost functions are fixed, and the one-step risk mappings have the same functional form across time). If $L_{t}(x)$ is decreasing in $x_{i}$ for each $t$ and some fixed $i$, then a control limit optimal policy in the $i$-th dimension exists, and the optimal control limit $\overline{x}_{t,i}(x_{t,-i})$ is increasing in $t$ for any fixed $x_{t,-i}$.
  • Monotonicity in risk-aversion level. Consider two time-homogeneous risk-averse optimal stopping instances with identical transition kernels and cost functions, and one-step risk mappings $\{\rho_{t}^{1}\}$ and $\{\rho_{t}^{2}\}$, respectively. Let the corresponding value functions be $v_{t}^{1}$ and $v_{t}^{2}$. For some fixed $i$, suppose a control limit optimal policy in the $i$-th dimension exists in both instances, with control limit functions denoted by $\overline{x}_{t,i}^{1}$ and $\overline{x}_{t,i}^{2}$, respectively. If the first instance is more risk-averse, i.e., $\rho_{t}^{1}(Z)\geq\rho_{t}^{2}(Z)$ for each $t$ and $Z\in\mathcal{Z}_{t+1}$, then $v_{t}^{1}\geq v_{t}^{2}$ and $\overline{x}_{t,i}^{1}(x_{t,-i})\geq\overline{x}_{t,i}^{2}(x_{t,-i})$ for each $t$ and $x_{t,-i}$.
@@
The third statement formalizes the effect of increased risk aversion under identical dynamics and costs. Its proof is a straightforward backward-induction argument and is omitted. For illustration, consider one-step mean–CVaR mappings $\rho_{t}^{i}(\cdot)=(1-\alpha_{i}){\mathbb{E}}(\cdot|\mathcal{F}_{t})+\alpha_{i}\operatorname*{CVaR}_{\gamma,t}(\cdot)$ (with fixed $\gamma\in[0,1]$) or pure CVaR mappings $\rho_{t}^{i}(\cdot)=\operatorname*{CVaR}_{\alpha_{i},t}(\cdot)$ for $i=1,2$. If $0\leq\alpha_{2}\leq\alpha_{1}\leq 1$, then $\rho_{t}^{1}(Z)\geq\rho_{t}^{2}(Z),~\forall Z\in\mathcal{Z}_{t+1}$. Intuitively, $\rho_{t}^{1}$ places more weight on the tail or on a more extreme tail and is therefore more risk-averse. Consequently, $v_{t}^{1}\geq v_{t}^{2}$, and the more risk-averse instance tends to stop earlier.